A rectangular lot has a surface of 375m^2. The length of one side is 60% of the other side. Determine the measure of each side of the lot.
Answer by Alan3354(69443)

A rectangular lot has a surface of 375m^2. The length of one side is 60% of the other side. Determine the measure of each side of the lot.
----------
x = long side
Area = x*0.6x = 375
x^2 = 625
x = 25 meters
--> 15 by 25